OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos 2x 2 sin2x - 2 sin x cos x + 1 2 sin x 2 sin2x + 2 sin x cos x - 1 2 sin2x - 2 sin x cos x - 1

OpenStudy (anonymous):

Im confused @mustafa2014

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

\[2 \sin2x + 2 \sin x \cos x - 1\] or \[sin^2x + 2 sin x \cos x - 1\]

OpenStudy (anonymous):

\[2\sin^2x + 2 \sin x \cos x - 1 \]

OpenStudy (anonymous):

@mondona

OpenStudy (anonymous):

ohh i think i understand

OpenStudy (anonymous):

\[\sin 2x=2 sin(x) cos(x)\] \[\cos 2x=cos^2x-sin^2x\] \[\sin 2x-cos 2x=2 sin(x) cos(x)-(cos^2x-sin^2x)\] \[\sin 2x-cos 2x=2 sin(x) cos(x)-cos^2x+sin^2x\] * \[\sin^2x+cos^2x=1\] \[\cos^2x=1-sin^2x\]* \[\sin 2x-cos 2x=2 sin(x) cos(x)-(1-sin^2x)+sin^2x\] \[\sin 2x-cos 2x=2 sin(x) cos(x)-1+sin^2x+sin^2x\] \[\sin 2x-cos 2x=2 sin(x) cos(x)-1+2sin^2x\]

OpenStudy (anonymous):

so it could be C

OpenStudy (anonymous):

@mustafa2014

OpenStudy (anonymous):

yes :)